\begin{references}
\bibitem{gronau1}  M. Gronau, Phys. Rev. {\bf D58}, 037301 (1998)

\bibitem{gronau2}  M. Gronau, Phys. Lett. {\bf B256}, 172 (1991)

\bibitem{xing}  Z. Z. Xing, Phys. Rev. D, hep-ph/9804434

\bibitem{gronau3}  M. Gronau, Phys. Rev. D, hep-ph/9807447

\bibitem{venezeno}  G. Veneziano, Nuovo Cimento {\bf 57A}, 190 (1960)

\bibitem{donoghue}  J. F. Donoghue et al. Phys. Rev. Lett. {\bf 77}, 2178
(1996)

\bibitem{falk}  A. F. Falk et al. Phys. Rev. {\bf D57}, 4290 (1998)

\bibitem{caprini}  I. Caprini, L. Micu and C. Bourrely, Phys. Rev. {\bf D60}%
, 074016 (1999)

\bibitem{neubert}  See for instance, H. Neubert and B. Stech,
hep-ph/9705292: To appear in the second edition of Heavy Flavors, edited by
A. J. Buras and M. Linder (World Scientific, Singapore). The refrences to
original literature can be found in this review).

\bibitem{wolfenstein}  L. Wolfenstein, Phys. Rev. Lett. {\bf 51}, 1945
(1983); N. Cabbibo, Phys. Rev. Lett. {\bf 10}, 531 (1963); M. Kobayski and
M. Maskawa, Prog. Theor. Phys. {\bf 49}, 652 (1973)
\end{references}

\documentstyle[aps]{revtex}

\begin{document}
\title{The decays $\bar{B}\rightarrow \bar{K}D$ and $\bar{B}\rightarrow \bar{K}\bar{%
D}$ and final state interactions}
\author{Fayyazuddin}
\address{National Center for Physics, Quaid-i-Azam University, Islamabad 45320,\\
Pakistan.}
\maketitle

\begin{abstract}
The decays $\bar{B}\rightarrow \bar{K}D$ and $\bar{B}\rightarrow \bar{K}\bar{%
D}$ taking into account final state interactions are discussed. These decays
are described by four strong phases $\delta _0,\delta _1,\bar{\delta}_0,\bar{%
\delta}_1$ (subscripts 0 and 1 refers to $I=0$ and $I=1$ final states), one
weak phase $\gamma $ and four real amplitudes. It is argued that strong
interaction dynamics implies $\bar{\delta}_1=0,\delta _0=-\delta _1$.
Rescattering has significant effects on weak amplitudes. Taking into
account, rescattering, we find that direct CP--violating asymmetry in these
decays may lie in the range $\mp 0.023\sin \gamma \leq {\cal A}_{1,2}\leq
\mp 0.086\sin \gamma $.
\end{abstract}

The weak decays $\bar{B}\rightarrow \bar{K}D$ and $\bar{B}\rightarrow \bar{K}%
\bar{D}$ taking into account final state interactions have been studied by
several authors \cite{gronau1,gronau2,xing,gronau3}. These decays are
described by four real amplitudes, four strong phases $\delta _0,\delta _1,%
\bar{\delta}_0,\bar{\delta}_1$ and one weak phase $\gamma $. The effective
Lagrangian which describes these decays are given by 
\begin{mathletters}
\label{01}
\begin{eqnarray}
L_{\text{eff}} &=&\frac{G_F}{\sqrt{2}}V_{cb}V_{us}^{*}\left[ \bar{s}\gamma
^\mu \left( 1-\gamma _5\right) u\right] \left[ \bar{c}\gamma _\mu \left(
1-\gamma _5\right) b\right]  \label{01a} \\
L_{\text{eff}} &=&\frac{G_F}{\sqrt{2}}V_{ub}V_{cs}^{*}\left[ \bar{s}\gamma
^\mu \left( 1-\gamma _5\right) c\right] \left[ \bar{u}\gamma _\mu \left(
1-\gamma _5\right) b\right]  \label{01b}
\end{eqnarray}
Since the effective weak Lagrangians for these decays have $\Delta I=1/2$,
the isospin analysis \cite{gronau3} for $\bar{K}D:$%
\end{mathletters}
\begin{mathletters}
\label{02}
\begin{eqnarray}
A\left( B^{-}\rightarrow K^{-}D^0\right) &=&2f_1e^{i\delta _1}  \label{02a}
\\
A\left( \bar{B}^0\rightarrow K^{-}D^{+}\right) &=&\left[ f_1e^{i\delta
_1}+f_0e^{i\delta _0}\right]  \label{02b} \\
A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) &=&\left[ f_1e^{i\delta
_1}-f_0e^{i\delta _0}\right]  \label{02c}
\end{eqnarray}
where $\delta _0$ and $\delta _1$ are the phase shifts for $I=0$ and $I=1$
isospin states. On the other hand for $\bar{K}\bar{D}$ we have 
\end{mathletters}
\begin{mathletters}
\label{03}
\begin{eqnarray}
A\left( \bar{B}^0\rightarrow \bar{K}^0\bar{D}^0\right) &=&2\bar{f}%
_1e^{i\gamma }e^{i\bar{\delta}_1}  \label{03a} \\
A\left( B^{-}\rightarrow K^{-}\bar{D}^0\right) &=&e^{i\gamma }\left[ \bar{f}%
_1e^{i\bar{\delta}_1}+\bar{f}_0e^{i\bar{\delta}_0}\right]  \label{03b} \\
A\left( B^{-}\rightarrow \bar{K}^0D^{-}\right) &=&e^{i\gamma }\left[ -\bar{f}%
_1e^{i\bar{\delta}_1}+\bar{f}_0e^{i\bar{\delta}_0}\right]  \label{03c}
\end{eqnarray}
In addition to above decays, other decays relevant to our analysis are 
\end{mathletters}
\begin{eqnarray}
A\left( \bar{B}^0\rightarrow \pi ^{+}D_s^{-}\right) &=&\bar{f}e^{i\gamma
}e^{i\bar{\delta}}=\sqrt{2}A\left( B^{-}\rightarrow \pi ^0D_s^{-}\right)
\label{04} \\
A\left( B^{-}\rightarrow \eta D_s^{-}\right) &=&\frac 1{\sqrt{6}}e^{i\gamma
}\left[ \bar{f}e^{i\bar{\delta}}-2\left( -\bar{f}_1e^{i\bar{\delta}}+\bar{f}%
_0e^{i\bar{\delta}_0}\right) \right]  \label{05}
\end{eqnarray}
where in writting Eq. (\ref{05}), we have used SU(3) which gives 
\begin{equation}
\sqrt{6}A\left( B^{-}\rightarrow \eta D_s^{-}\right) =A\left( \bar{B}%
^0\rightarrow \pi ^{+}D_s^{-}\right) -2A\left( B^{-}\rightarrow \bar{K}^0%
\bar{D}\right)  \label{06}
\end{equation}

In this paper, we address two questions: Is it possible to get some
constraints on strong phases ? How the weak amplitudes are affected by
rescattering ? In answer to first question, our analysis implies 
\begin{equation}
\delta _0=-\delta _1,\,\,\,\,\,\,\bar{\delta}_1=0.  \label{07}
\end{equation}
We find that rescattering has significant effect on some observables for
these decays.

Let us consider the scattering processes: 
\begin{eqnarray*}
\bar{K}+D &\rightarrow &\bar{K}+D \\
\bar{K}+\bar{D} &\rightarrow &\bar{K}+\bar{D}
\end{eqnarray*}
where $\bar{K}\equiv \left( \bar{K}^0,\,K^{-}\right) ,\,D\equiv \left(
D^{+},\,D^0\right) ,\,\bar{D}\equiv \left( \bar{D}^0,\,D^{-}\right) $. We
note that in $s$ or $u$-channel only the states with quark structure $s\bar{c%
}$ can be exchanged ($s,u$ and $t$ are Mandelstam variables). An important
consequence of this is that $s$-channel is exotic for $\bar{K}D$ scattering
whereas $u$-channel is exotic for $\bar{K}\bar{D}$ scattering. Since $s\bar{c%
}$ state has isospin $I=0$, this exchange in $s$-channel gives isospin
projection operator $-2P_0$ where as $I=0$ exchange in $u$-channel gives
isospin projection operator $P_1-P_0$. In $t$-channel, $I=1\left( \rho
\right) $ exchange gives an isospin factor $P_1-3P_0$ where as $I=0\left(
\omega \right) $ exchange gives an isospin factor $P_1+P_0$. Assuming $\rho
-\omega $ degeneracy, $I=1$ and $I=0$ exchanges give $\left( P_1-3P_0\right)
+\left( P_1+P_0\right) =2\left( P_1-P_0\right) $ isospin projection operator
for $\bar{K}D$ scattering where as for $\bar{K}\bar{D}$ scattering. we get $%
\left( P_1-3P_0\right) -\left( P_1+P_0\right) =-4P_0$. Hence for $\bar{K}D$
and $\bar{K}\bar{D}$, the scattering amplitudes can be written as 
\begin{eqnarray}
M &=&-gA\left( P_1-P_0\right)  \label{08} \\
\bar{M} &=&-g\bar{A}\left( -2P_0\right)  \label{09}
\end{eqnarray}
Eqs. (\ref{08}) and (\ref{09}) imply $\delta _0=-\delta _1,\bar{\delta}_1=0$
viz Eq. (\ref{07}).

It is clear from Eq. (\ref{08}) that scattering matrices for $\bar{K}D$ for $%
Q=-1$ and $Q=0$ states are given by respectively 
\begin{equation}
M_{-0}=-gA,\,\,\,\,\,\,\left( K^{-}D^0\rightarrow K^{-}D^0\right) 
\label{10}
\end{equation}
\begin{equation}
M=
\begin{tabular}{c|cc}
& $00$ & $-+$ \\ \hline
$00$ & $0$ & $-gA$ \\ 
$-+$ & $-gA$ & $0$%
\end{tabular}
\label{11}
\end{equation}
From Eq. (\ref{08}), we get for Q=0 and Q=-1 states of $\bar{K}\bar{D}$, the
scattering matrices 
\begin{eqnarray}
\bar{M}_{00} &=&0,\,\,\,\,\,\,\,\,\left( \bar{K}^0\bar{D}^0\rightarrow \bar{K%
}^0\bar{D}^0\right)   \label{12} \\
\bar{M} &=&
\begin{tabular}{ccc}
& $0-$ & $-0$ \\ 
$0-$ & $g\bar{A}$ & $-g\bar{A}$ \\ 
$-0$ & $-g\bar{A}$ & $g\bar{A}$%
\end{tabular}
\label{13}
\end{eqnarray}
However, we note that for $\bar{K}\bar{D}$ scattering other channels $\pi
^{+}D_s^{-},\pi ^0D_s^{-}$ and $\eta D_s^{-}$ are also open. Hence we must
extend the scattering matrices given in Eqs. (\ref{12}) and (\ref{13}) to
take into account the inelastic channels. Using SU(3), we get the scattering
matrices 
\begin{equation}
\bar{M}\left( Q=0\right) =
\begin{tabular}{c|cc}
& $00$ & $\pi ^{+}s^{-}$ \\ \hline
$00$ & $0$ & $-g{A}$ \\ 
$\pi ^{+}s^{-}$ & $-g{A}$ & $0$%
\end{tabular}
\label{14}
\end{equation}
and 
\begin{equation}
\bar{M}\left( Q=-1\right) =
\begin{tabular}{ccccc}
& $0-$ & $-0$ & $\pi ^0s^{-}$ & $\eta s^{-}$ \\ 
$0-$ & $g\bar{A}$ & $-g\bar{A}$ & $\frac 1{\sqrt{2}}gA$ & $\frac 1{\sqrt{6}}%
g\left( 2\bar{A}-A\right) $ \\ 
$-0$ & $-g\bar{A}$ & $g\bar{A}$ & $-\frac 1{\sqrt{2}}gA$ & $\frac 1{\sqrt{6}}%
g\left( 2\bar{A}-A\right) $ \\ 
$\pi ^0s^{-}$ & $\frac 1{\sqrt{2}}gA$ & $-\frac 1{\sqrt{2}}gA$ & 0 & 0 \\ 
$\eta s^{-}$ & $\frac 1{\sqrt{6}}g\left( 2\bar{A}-A\right) $ & $\frac 1{%
\sqrt{6}}g\left( 2\bar{A}-A\right) $ & 0 & $-\frac 23g\left( A+\bar{A}%
\right) $%
\end{tabular}
\label{15}
\end{equation}
To proceed further we note that in terms of Regge phenomenology, exotic $u$%
-channel implies exchange degeneracy i.e. in $t$-channel $\rho -A_2$ as well
as $\omega -f$ trajectories are exchange degenrate. Taking this into
account, it is convenient to express amplitudes $A$ and $\bar{A}$ in
Veneziano representation \cite{venezeno} 
\begin{eqnarray}
A &=&\left[ \frac{\Gamma \left( 1-\alpha _{D_s^{*}}\left( u\right) \right)
\Gamma \left( 1-\alpha \left( t\right) \right) }{\Gamma \left( 1-\alpha
_{D_s^{*}}\left( u\right) -\alpha \left( t\right) \right) }\right] 
\label{16} \\
\bar{A} &=&\left[ \frac{\Gamma \left( 1-\alpha _{D_s^{*}}\left( s\right)
\right) \Gamma \left( 1-\alpha \left( t\right) \right) }{\Gamma \left(
1-\alpha _{D_s^{*}}\left( s\right) -\alpha \left( t\right) \right) }\right] 
\label{17}
\end{eqnarray}
We will take linear Regge trajectories viz 
\begin{eqnarray}
\alpha  &=&\alpha _0+\alpha ^{\prime }t  \nonumber \\
\alpha _{D_s^{*}}\left( u\right)  &=&\alpha _{D_s^{*}}\left( 0\right)
+\alpha ^{\prime }u  \nonumber \\
\alpha _{D_s^{*}}\left( s\right)  &=&\alpha _{D_s^{*}}\left( 0\right)
+\alpha ^{\prime }s  \label{18}
\end{eqnarray}
We assume universal slope viz 
\begin{equation}
\alpha ^{\prime }\simeq 0.94\text{ GeV}^{-2}\approx \frac 1{s_0}  \label{19}
\end{equation}
and take $\alpha _0=0.46$. In actual numerical evaluation we will put $%
\alpha _0=1/2$ and $s_0=1$ GeV$^2$. Note that $\alpha _{D_s^{*}}\left(
0\right) =1-\alpha ^{\prime }m_{D_s^{*}}^2$ for large $s$, we get from Eqs. (%
\ref{16}) and (\ref{17})$:$%
\begin{eqnarray}
A &\rightarrow &\frac \pi {\sin \pi \alpha \left( t\right) \Gamma \left(
\alpha \left( t\right) \right) }\left( \frac s{s_0}\right) ^{\alpha \left(
t\right) }  \label{20} \\
\bar{A} &\rightarrow &\frac{\pi e^{-i\pi \alpha \left( t\right) }}{\sin \pi
\alpha \left( t\right) \Gamma \left( \alpha \left( t\right) \right) }\left( 
\frac s{s_0}\right) ^{\alpha \left( t\right) }  \label{21}
\end{eqnarray}
From Eqs. (\ref{17}), we note that $1-\alpha _{D_s^{*}}\left( s\right) =0$
gives a pole at $s=m_{D_s^{*}}^2$ and $1-\alpha \left( t\right) =0$ gives a
pole at $t=m_\rho ^2$. Using this property of Venezeno representation, we
get 
\begin{equation}
-g=2g_{D_s^{*}DK}^2=4g_{\rho K\bar{K}}g_{\rho D\bar{D}}  \label{22}
\end{equation}
Using the usual parameterization for $g_{D_s^{*}DK}^2$, we get 
\begin{equation}
g_{D_s^{*}DK}^2=\gamma _D^2\frac{m_{D_s^{*}}^2}{f_K^2}\simeq 50  \label{23}
\end{equation}
for $\gamma _D=1/2$.

We now come to the second question viz the effect of rescattering on the
weak amplitudes. First we discuss the rescattering for $\bar{K}D$ system.
Two particle unitarity gives \cite{donoghue,falk,caprini} 
\begin{eqnarray}
Disc\,A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) &=&\frac 1{32\pi }%
\frac{\left| {\bf p}\right| }s\int d\Omega A\left( \bar{B}^0\rightarrow
K^{-}D^{+}\right) M^{*}\left( K^{-}D^{+}\rightarrow \bar{K}^0D^0\right) 
\nonumber \\
&=&\frac 1{32\pi }\frac 1{\sqrt{s}\left| {\bf p}\right| }\int_{-2\left| {\bf %
p}\right| ^2}^0dt\,A\left( \bar{B}^0\rightarrow K^{-}D^{+}\right)
M^{*}\left( K^{-}D^{+}\rightarrow \bar{K}^0D^0\right)  \label{24}
\end{eqnarray}
Using Eqs. (\ref{11}) and (\ref{20}), we get 
\begin{equation}
Disc\,A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) =\frac 1{32\pi }\frac 
1{\sqrt{s}\left| {\bf p}\right| }\frac{\left( -g\right) \pi A\left( \bar{B}%
^0\rightarrow K^{-}D^{+}\right) }{\Gamma \left( \alpha \left( 0\right)
\right) \sin \pi \alpha \left( 0\right) }\int_{-2\left| {\bf p}\right|
^2}^0e^{\alpha \left( t\right) \ln \left( s/s_0\right) }dt  \label{25}
\end{equation}
where we have put \cite{falk}, $\alpha \left( t\right) \approx \alpha \left(
0\right) $ in $\Gamma \left( \alpha \left( t\right) \right) $ and $\sin \pi
\alpha \left( t\right) $. Hence we get, taking $\alpha \left( 0\right)
\simeq 1/2$ and using linear Regge trajectory: 
\begin{equation}
Disc\,A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) \approx -\frac g{16%
\sqrt{\pi }}\frac{\left( s/s_0\right) ^{\alpha _0-1}}{\ln \left(
s/s_0\right) }A\left( \bar{B}^0\rightarrow K^{-}D^{+}\right)  \label{26}
\end{equation}
We now use dispersion relation to obtain $A\left( \bar{B}^0\rightarrow
K^{-}D^{+}\right) $ \cite{donoghue,falk}: 
\begin{eqnarray}
A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) _{\text{FSI}} &=&-\frac g{16%
\sqrt{\pi }}\frac 1{\ln \left( m_B^2/s_0\right) }A\left( \bar{B}%
^0\rightarrow K^{-}D^{+}\right) \times \frac 1\pi \int_{\left(
m_D+m_K\right) ^2}^\infty \frac{\left( s/s_0\right) ^{\alpha _0-1}}{s-m_B^2}
\nonumber \\
&=&-\frac g{16\sqrt{\pi }}\frac{\sqrt{s_0}}{m_B}\frac 1{\ln \left(
m_B^2/s_0\right) }\frac 1\pi \left[ i\pi +\ln \frac{1+x}{1-x}\right] A\left( 
\bar{B}^0\rightarrow K^{-}D^{+}\right)  \nonumber \\
&\equiv &\epsilon e^{i\theta }A\left( \bar{B}^0\rightarrow K^{-}D^{+}\right)
\label{27}
\end{eqnarray}
where in evaluating the dispersion integral we have replaced \cite{falk} $%
\ln \left( s/s_0\right) $ by $\ln \left( m_B^2/s_0\right) $. In Eq. (\ref{27}%
) $x,\epsilon $ and $\theta $ are given by 
\begin{eqnarray}
x &=&\frac{m_D+m_K}{m_B}\simeq 0.447  \nonumber \\
\epsilon &=&-\frac g{16\sqrt{\pi }}\frac{s_0}{m_B\ln \left( m_B^2/s_0\right) 
}\sqrt{1+\frac 1{\pi ^2}\left( \ln \frac{1+x}{1-x}\right) ^2}  \nonumber \\
&=&-g\left( 2.01\times 10^{-3}\right) =0.20  \nonumber \\
\theta &=&\tan ^{-1}\left[ \frac \pi {\ln \frac{1+x}{1-x}}\right] \approx
73^0  \label{28}
\end{eqnarray}
where we have used $-g\approx 100$ as given in Eqs. (\ref{22}) and (\ref{23})

Similarly, we obtain 
\begin{equation}
A\left( \bar{B}^0\rightarrow K^{-}D^{+}\right) _{\text{FSI}}=\epsilon
e^{i\theta }A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right)  \label{30}
\end{equation}
For the decays $\bar{B}\rightarrow \bar{K}\bar{D}$, we note that except for
the decays $\bar{B}^0\rightarrow \pi ^{+}D_s^{-}$, $B^{-}\rightarrow \pi
^0D_s^{-}$, $B^{-}\rightarrow \eta D_s^{-}$, all other decays are either
color suppressed or are given by annihilation amplitude. Thus in evaluating
rescattering correction, we retain only the dominent amplitudes. From Eqs. (%
\ref{14}) and (\ref{15}), following the same procedure as above, we get 
\begin{eqnarray}
A\left( \bar{B}^0\rightarrow \bar{K}^0\bar{D}^0\right) _{\text{FSI}}
&=&\epsilon e^{i\theta }A\left( \bar{B}^0\rightarrow \pi ^{+}D_s^{-}\right)
\label{31} \\
A\left( B^{-}\rightarrow \bar{K}^0D^{-}\right) _{\text{FSI,}\pi } &=&\frac 12%
\epsilon e^{i\theta }A\left( \bar{B}^0\rightarrow \pi ^{+}D_s^{-}\right) 
\nonumber \\
&=&-A\left( B^{-}\rightarrow K^{-}\bar{D}^0\right) _{\text{FSI,}\pi }
\label{32} \\
A\left( B^{-}\rightarrow \bar{K}^0D^{-}\right) _{\text{FSI,}\eta } &=&\frac 1%
{\sqrt{6}}\epsilon e^{i\theta }\left[ 1-2\frac{\ln \left( m_B^2/s_0\right) }{%
\sqrt{\pi ^2+\left( \ln \left( m_B^2/s_0\right) \right) ^2}}e^{i\pi
/2}e^{-i\chi }\right] A\left( \bar{B}^0\rightarrow \eta D_s^{-}\right) 
\nonumber \\
&=&A\left( B^{-}\rightarrow K^{-}\bar{D}^0\right) _{\text{FSI}}  \label{33}
\end{eqnarray}
where 
\begin{equation}
\chi =\tan ^{-1}\frac \pi {\ln \left( m_B^2/s_0\right) }\approx \tan
^{-1}\left( 0.94\right) \approx \frac \pi 4  \label{34}
\end{equation}
In order to simplify the calculation, we replace 
\begin{equation}
\frac{\ln \left( m_B^2/s_0\right) }{\sqrt{\left( \ln \left( m_B^2/s_0\right)
\right) ^2+\pi ^2}}=0.727\approx \frac 1{\sqrt{2}}  \label{35}
\end{equation}
Hence, from Eq. (\ref{33}), using Eqs. (\ref{34}) and (\ref{35}), we get 
\begin{eqnarray}
A\left( B^{-}\rightarrow \bar{K}^0D^{-}\right) _{\text{FSI,}\eta } &=&-\frac 
i{\sqrt{6}}\epsilon e^{i\theta }A\left( \bar{B}^0\rightarrow \eta
D_s^{-}\right)  \nonumber \\
&=&A\left( B^{-}\rightarrow K^{-}\bar{D}^0\right) _{\text{FSI,}\eta }
\label{36}
\end{eqnarray}
Now using Eq. (\ref{06}), and neglecting the contribution of $A\left(
B^{-}\rightarrow \bar{K}^0D^{-}\right) $ compared to $A\left( \bar{B}%
^0\rightarrow \pi ^{+}D_s^{-}\right) $, we obtain 
\begin{eqnarray}
A\left( B^{-}\rightarrow \bar{K}^0D^{-}\right) _{\text{FSI,}\pi ^0-\eta } &=&%
\frac 12\epsilon e^{i\theta }\left( 1-\frac i3\right) A\left( \bar{B}%
^0\rightarrow \pi ^{+}D_s^{-}\right)  \nonumber \\
&=&\frac{\sqrt{10}}6\epsilon e^{i\left( \theta -\phi \right) }A\left( \bar{B}%
^0\rightarrow \pi ^{+}D_s^{-}\right)  \label{37} \\
A\left( B^{-}\rightarrow K^{-}\bar{D}^0\right) _{\text{FSI,}\pi ^0-\eta }
&=&-\frac 12\epsilon e^{i\theta }\left( 1+\frac i3\right) A\left( \bar{B}%
^0\rightarrow \pi ^{+}D_s^{-}\right)  \nonumber \\
&=&-\frac{\sqrt{10}}6\epsilon e^{i\left( \theta +\phi \right) }A\left( \bar{B%
}^0\rightarrow \pi ^{+}D_s^{-}\right)  \label{38}
\end{eqnarray}
where 
\begin{eqnarray}
\phi &=&\tan ^{-1}\left( \frac 13\right) \approx 18^0  \nonumber \\
\theta +\phi &=&91^0\approx 90^0  \nonumber \\
\theta -\phi &=&55^0  \label{39}
\end{eqnarray}
Taking into rescattering, Eqs. (\ref{02}) and (\ref{03}) are modified 
\begin{mathletters}
\label{40}
\begin{eqnarray}
A\left( B^{-}\rightarrow K^{-}D^0\right) &=&2f_1\left( 1+\epsilon e^{i\theta
}\right) e^{i\delta _1}  \label{40a} \\
A\left( \bar{B}^0\rightarrow K^{-}D^{+}\right) &=&f_1\left( 1+\epsilon
e^{i\theta }\right) e^{i\delta _1}+f_0\left( 1-\epsilon e^{i\theta }\right)
e^{i\delta _0}  \label{40b} \\
A\left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) &=&f_1\left( 1+\epsilon
e^{i\theta }\right) e^{i\delta _1}-f_0\left( 1-\epsilon e^{i\theta }\right)
e^{i\delta _0}  \label{40c}
\end{eqnarray}
\end{mathletters}
\begin{mathletters}
\label{41}
\begin{eqnarray}
A\left( \bar{B}^0\rightarrow \bar{K}^0\bar{D}^0\right) &=&e^{i\gamma }\left[
2\bar{f}_1+\epsilon \bar{f}e^{i\left( \theta +\bar{\delta}\right) }\right]
\label{41a} \\
A\left( B^{-}\rightarrow K^{-}\bar{D}^0\right) &=&e^{i\gamma }\left[ \left( 
\bar{f}_1+\bar{f}_0e^{i\bar{\delta}_0}\right) -\frac{\sqrt{10}}6\epsilon
e^{i\left( \theta +\phi +\bar{\delta}\right) }\bar{f}\right]  \label{41b} \\
A\left( B^{-}\rightarrow \bar{K}^0D^{-}\right) &=&e^{i\gamma }\left[ \left( -%
\bar{f}_1+\bar{f}_0e^{i\bar{\delta}_0}\right) +\frac{\sqrt{10}}6\epsilon
e^{i\left( \theta -\phi +\bar{\delta}\right) }\bar{f}\right]  \label{41c}
\end{eqnarray}
The observables which are significantly affected by rescattering can be
easily obtained from Eqs. (\ref{40}) and (\ref{41}). From these equations,
we get 
\end{mathletters}
\begin{eqnarray}
R &=&\frac{\Gamma \left( \bar{B}^0\rightarrow \bar{K}^0\bar{D}^0\right) }{%
\Gamma \left( \bar{B}^0\rightarrow \bar{K}^0D^0\right) }  \nonumber \\
&\simeq &\frac{4\bar{f}_1^2\left[ 1+\epsilon \frac{\bar{f}}{\bar{f}_1}%
+\epsilon ^2\frac{\bar{f}^2}{4\bar{f}_1^2}\right] }{f_1^2+f_0^2-2f_1f_0\cos
\left( \delta _1-\delta _0\right) }  \label{42}
\end{eqnarray}
(where in the denominator, we have neglected the terms containing $\epsilon $%
). 
\begin{eqnarray}
R_{1,2} &\equiv &\frac{\Gamma \left( B^{-}\rightarrow K^{-}D_{1,2}\right)
+\Gamma \left( B^{+}\rightarrow K^{+}D_{1,2}\right) }{\Gamma \left(
B^{-}\rightarrow K^{-}D^0\right) }  \nonumber \\
&=&1+\frac 14\left[ r_1^2+r_0^2+2r_1r_0\cos \bar{\delta}_0\right] \mp \cos
\gamma \left[ r_1\cos \delta _1+r_0\cos \left( \delta _1-\bar{\delta}%
_0\right) \right] -\frac{\sqrt{10}}6\epsilon r\cos \left( \theta +\phi +\bar{%
\delta}-\delta _1\right)  \label{43}
\end{eqnarray}
\begin{eqnarray}
{\cal A}_{1,2} &\equiv &\frac{\Gamma \left( B^{-}\rightarrow
K^{-}D_{1,2}\right) -\Gamma \left( B^{+}\rightarrow K^{+}D_{1,2}\right) }{%
\Gamma \left( B^{-}\rightarrow K^{-}D^0\right) }  \nonumber \\
&=&\mp \sin \gamma \left[ r_1\sin \delta _1+r_0\sin \left( \delta _1-\bar{%
\delta}_0\right) +\frac{\sqrt{10}}6\epsilon r\sin \left( \theta +\phi +\bar{%
\delta}-\delta _1\right) \right]  \label{44}
\end{eqnarray}
where $D_{1,2}=\left( D^0\mp \bar{D}^0\right) /\sqrt{2}$ are CP-eigenstates
with CP $=+1,-1$ respectively. $r_1,r_0$ and $r$ are given by 
\begin{equation}
r_1=\frac{\bar{f}_1}{f_1},\text{ }r_0=\frac{\bar{f}_0}{f_1}\text{ and }r=%
\frac{\bar{f}}{f_1}  \label{45}
\end{equation}

So far our analysis is general. To proceed further, we note that \cite
{neubert} that these decays are determined by the tree amplitudes $T\left( 
\bar{T}\right) $, the color suppressed amplitudes $C\left( \bar{C}\right) $
and annihilation amplitude $\bar{A}$. In terms of these amplitudes 
\begin{eqnarray}
f_1 &=&\frac{G_F}{\sqrt{2}}\left| V_{cb}V_{us}^{*}\right| \frac 12\left(
T+C\right)  \label{46} \\
f_0 &=&\frac{G_F}{\sqrt{2}}\left| V_{cb}V_{us}^{*}\right| \frac 12\left(
T-C\right)  \label{47} \\
\bar{f}_1 &=&\frac{G_F}{\sqrt{2}}\left| V_{ub}V_{cs}^{*}\right| \frac 12\bar{%
C}  \label{48} \\
\bar{f}_0 &=&\frac{G_F}{\sqrt{2}}\left| V_{ub}V_{cs}^{*}\right| \frac 12%
\left( \bar{C}+2\bar{A}\right)  \label{49} \\
\bar{f} &=&\frac{G_F}{\sqrt{2}}\left| V_{ub}V_{cs}^{*}\right| \bar{T}
\label{50}
\end{eqnarray}
Note that in the Wolfenstein representation of CKM matrix \cite{wolfenstein}%
: 
\begin{equation}
\frac{\left| V_{ub}V_{cs}^{*}\right| }{\left| V_{cb}V_{us}^{*}\right| }=%
\sqrt{\rho ^2+\eta ^2}  \label{51}
\end{equation}
In the factorization anastz, these amplitudes are given by \cite{neubert} 
\begin{eqnarray}
T &=&a_1f_KF_0^{B-D}\left( m_K^2\right) \left( m_B^2-m_D^2\right)  \nonumber
\\
C &=&a_2f_DF_0^{B-K}\left( m_D^2\right) \left( m_B^2-m_K^2\right) =\bar{C} 
\nonumber \\
\bar{T} &=&a_1f_{D_s}F_0^{B-\pi }\left( m_{D_s}^2\right) \left( m_B^2-m_\pi
^2\right)  \nonumber \\
\bar{A} &=&a_1f_BF_0^{D-K}\left( m_B^2\right) \left( m_D^2-m_K^2\right)
\label{52}
\end{eqnarray}
where $F_0\left( t\right) $ is scalar form factor for $B$ to $P$ transition $%
\left( P=D,K\text{ or }\pi \right) $. For these form factors, we use the
following values 
\begin{equation}
F_0^{B-K}\left( m_D^2\right) \approx F_0^{B-\pi }\left( m_{D_s}^2\right)
\simeq 0.22  \label{53}
\end{equation}
and \cite{neubert} 
\begin{equation}
F_0^{B-D}\left( m_K^2\right) \approx 0.587  \label{54}
\end{equation}
Using \cite{neubert}, $f_D\simeq 200$ MeV, $f_{D_s}\simeq 240$ MeV, $%
f_B\simeq 180$ MeV, $f_K=158$ MeV and $a_2/a_1\simeq 0.26$, we get 
\begin{eqnarray}
r_1 &=&\frac{\bar{f}_1}{f_1}=\sqrt{\rho ^2+\eta ^2}\frac{\bar{C}}{C+T}%
\approx 0.04  \nonumber \\
r_0 &=&\frac{\bar{f}_0}{f_1}=\sqrt{\rho ^2+\eta ^2}\frac{\bar{C}+2\bar{A}}{%
C+T}\approx 0.08  \nonumber \\
r &=&\frac{\bar{f}}{f_1}=\sqrt{\rho ^2+\eta ^2}\frac{\bar{T}}{C+T}\approx
0.41  \label{55}
\end{eqnarray}
where we have 
\begin{equation}
\sqrt{\rho ^2+\eta ^2}=0.36  \label{56}
\end{equation}

We now discuss the consequences of our main results given in Eqs. (\ref{42})
and (\ref{44}). First we note that it follows from Eq. (\ref{14}), that $%
\bar{\delta}$ can be put equal to zero (no elastic scattering for $\pi
D_s^{-}$). Hence, from Eq. (\ref{42}), we get

\begin{equation}
R=R_0\left[ 1+\epsilon \bar{r}\cos \theta +\frac{\epsilon ^2\bar{r}^2}4%
\right]  \label{57}
\end{equation}
where 
\begin{equation}
R_0=\frac{4\bar{f}_1^2}{f_1^2+f_0^2-2f_1f_0\cos \left( \delta _1-\delta
_0\right) }  \label{58}
\end{equation}
is the branching ratio in the absence of rescattering. From Eqs. (\ref{48}),
(\ref{50}) and (\ref{51}), we get 
\begin{equation}
\bar{r}\equiv \frac{\bar{f}}{\bar{f}_1}=\frac{2\bar{T}}{\bar{C}}\simeq 5.00
\label{59}
\end{equation}
Hence we obtain using $\epsilon =0.20$ and $\theta =73^0,$%
\begin{equation}
R=1.54R_0  \label{60}
\end{equation}
We now discuss the consequence of Eq. (\ref{44}) i.e. direct CP-violation in 
$\bar{B}$ decays.Let us first assume that final state interactions are taken
care of by the phases induced by rescattering. Hence we put $\delta _1$ and $%
\delta _1-\bar{\delta}_0$ equal to zero. Then, we get from Eq. (\ref{44}) 
\begin{eqnarray}
{\cal A}_{1,2} &=&\mp \sin \gamma \left[ \frac{\sqrt{10}}6\epsilon r\sin
\left( \theta +\phi \right) \right]  \nonumber \\
&=&\mp \sin \gamma \left[ \frac{\sqrt{10}}6\epsilon r\right]  \nonumber \\
&=&\mp \sin \gamma \left( 0.043\right)  \label{61}
\end{eqnarray}
since $\theta +\phi \approx 90^0.$

In general however 
\begin{equation}
{\cal A}_{1,2}=\mp \sin \gamma \left[ 0.04\sin \delta _1+0.08\sin \left(
\delta _1-\bar{\delta}_0\right) +0.04\sin \left( \frac \pi 2-\delta
_1\right) \right]  \label{62}
\end{equation}

But the structure of Eqs. (\ref{08}), (\ref{09}), (\ref{16}) and (\ref{17})
implies that $\bar{\delta}_0$ has the same sign as $\delta _0$, it is
therefore reasonable to conclude that $\delta _1-\bar{\delta}_0$ has the
same sign as $\delta _1-\delta _0$. Since $\delta _0=-\delta _1$, it follows
that (since $\delta _1$ has positive sign), that ${\cal A}_{1,2}$ should be
atleast $\mp \left( 0.043\right) \sin \gamma $. The phases $\delta _1,\bar{%
\delta}_0$ are expected to be small. As an example, let us take $\delta
_1=13^0,\delta _0=-13^0=\bar{\delta}_0$, then we get 
\begin{equation}
{\cal A}_{1,2}=\mp \sin \gamma \left( 0.086\right)  \label{63}
\end{equation}

The direct CP--violation is an important consequence of the standard model.
But in the absence of final state interactions, this parameter is zero. Our
analysis shows that even if strong phases $\delta $'s are negligible, the
rescattering gives a finite value for ${\cal A}_{1,2}$ which may be
experimentally measureable in future experiments. Even if our estimate of $%
\epsilon $ is off by a factor 2, ${\cal A}_{1,2}$ will still have at least
the value $\mp 0.023\sin \gamma $. Thus unless in an unlikely case that $%
\bar{\delta}_0$ has the same sign as $\delta _1$ and much greater than $%
\delta _1$, One may conclude that direct asymmetry parameter may lie in the
range 
\begin{equation}
\mp 0.023\sin \gamma \leq {\cal A}_{1,2}\leq \mp 0.086\sin \gamma  \label{64}
\end{equation}

\end{document}